OpenStudy (anonymous):

Simplify. Express the product as a rational exponent. x^1/3 times x^1/3

OpenStudy (anonymous):

i think it might be x^2/3

OpenStudy (zepp):

\[\large (x^\frac{1}{3})^2=x^\frac{2}{3}\]

OpenStudy (zepp):

Law of power of a power, you are right :)